# GaussianCube: Structuring Gaussian Splatting using Optimal Transport for 3D Generative Modeling

### v2

@@ Abstract
3D Gaussian Splatting (GS) have achieved considerable improvement over Neural Radiance Fields in terms of 3D fitting fidelity and rendering speed. However, this unstructured representation with scattered Gaussians poses a significant challenge for generative modeling. To address the problem, we introduce GaussianCube, a structured GS representation that is both powerful and efficient for generative modeling. We achieve this by first proposing a modified densification-constrained GS fitting algorithm which can yield high-quality fitting results using a fixed number of free Gaussians, and then re-arranging the Gaussians into a predefined voxel grid via Optimal Transport. The structured grid representation allows us to use standard 3D U-Net as our backbone in diffusion generative modeling without elaborate designs. Extensive experiments conducted on ShapeNet and OmniObject3D show that our model achieves state-of-the-art generation results both qualitatively and quantitatively, underscoring the potential of GaussianCube as a powerful and versatile 3D representation. Project page: https://gaussiancube.github.io/.
11footnotetext: Interns at Microsoft Research Asia.
@@ 1 Introduction
Recent advancements in generative modeling Ho et al. (2020); Goodfellow et al. (2020); Nichol and Dhariwal (2021); Dhariwal and Nichol (2021); Zhang et al. (2022); Karras et al. (2019) have led to significant progress in 3D content creation Wang et al. (2023); Müller et al. (2023); Cao et al. (2023); Tang et al. (2023c); Shue et al. (2023); Chan et al. (2022); Gao et al. (2022). Most of the prior works in this domain leverage variants of Neural Radiance Field (NeRF) Mildenhall et al. (2021) as their underlying 3D representations Chan et al. (2022); Tang et al. (2023c), which typically consist of an explicit and structured proxy representation and an implicit feature decoder. However, such hybrid NeRF variants have degraded representation power, particularly when used for generative modeling where a single implicit feature decoder is shared across all objects. Furthermore, the high computational complexity of volumetric rendering leads to both slow rendering speed and extensive memory costs. Recently, the emergence of 3D Gaussian Splatting (GS) Kerbl et al. (2023) has enabled high-quality reconstruction Xu et al. (2023); Luiten et al. (2023); Wu et al. (2023a) along with real-time rendering speed. The fully explicit characteristic of 3DGS also eliminates the need for a shared implicit decoder. Although 3DGS has been widely studied in scene reconstruction tasks, its spatially unstructured nature presents significant challenge when applying it to generative modeling.
@@
In this work, we introduce GaussianCube, a novel representation crafted to address the unstructured nature of 3DGS and unleash its potential for 3D generative modeling (see Table 1 for comparisons with prior works). Converting 3D Gaussians into a structured format without sacrificing their expressiveness is not a trivial task. We propose to first perform high-quality fitting using a fixed number of Gaussians and then organize them in a spatially structured manner. To keep the number of Gaussians fixed during fitting, a naive solution might omit the densification and pruning steps in GS, which, however, would significantly degrade the fitting quality. In contrast, we propose a densification-constrained fitting strategy, which retains the original pruning process yet constrains the number of Gaussians that perform densification, ensuring the total does not exceed a predefined maximum $N^{3}$ (32,768 in this paper). For the subsequent structuralization, we allocate the Gaussians across an $N\times N\times N$ voxel grid using Optimal Transport (OT). Consequently, our fitted Gaussians are systematically arranged within the voxel grid, with each grid containing a Gaussian feature. The proposed OT-based structuralization process achieves maximal spatial coherence, characterized by minimal total transport distances, while preserving the high expressiveness of the 3DGS.
Representation Spatially-structured Fully-explicit High-quality Reconstruction Efficient Rendering
Vanilla NeRF Mildenhall et al. (2021)
Neural Voxels Tang et al. (2023c)
Triplane Chan et al. (2022)
Gaussian Splatting Kerbl et al. (2023)
Our GaussianCube
Table 1: Comparison with prior 3D representations.
We perform 3D generative modeling with the proposed GaussianCube using diffusion models Ho et al. (2020). The spatially coherent structure of the Gaussians in our representation facilitates efficient feature extraction and permits the use of standard 3D convolutions to capture the correlations among neighboring Gaussians effectively. Therefore, we construct our diffusion model with standard 3D U-Net architecture without elaborate designs. It is worth noting that our diffusion model and the GaussianCube representation are generic, which facilitates both unconditional and conditional generation tasks.
@@
We conduct comprehensive experiments to verify the efficacy of our proposed approach. The model’s capability for unconditional generation is evaluated on the ShapeNet dataset Chang et al. (2015). Both the quantitative and qualitative comparisons indicate that our model surpasses all previous methods. Additionally, we perform class-conditioned generation on the OmniObject3D dataset Wu et al. (2023b), which is a extensive collection of real-world scanned objects with a broad vocabulary. Our model excels in producing semantically accurate 3D objects with complex geometries and realistic textures, outperforming the state-of-the-art methods. These experiments collectively demonstrate the strong capabilities of our GaussianCube and suggest its potential as a powerful and versatile 3D representation for a variety of applications. Some generated samples of our method is presented in Figure 1.
Figure 1: Samples of our generated 3D objects. Our model is able to create diverse objects with complex geometry and rich texture details.
@@ 2 Related Work
Radiance field representation. Radiance fields model ray interactions with scene surfaces and can be in either implicit or explicit forms. Early works of neural radiance fields (NeRFs) Mildenhall et al. (2021); Zhang et al. (2020); Park et al. (2021); Barron et al. (2022); Pumarola et al. (2021) are often in an implicit form, which represents scenes without defining geometry. These works optimize a continuous scene representation using volumetric ray-marching that leads to extremely high computational costs. Recent works introduce the use of explicit proxy representation followed by an implicit feature decoder to enable faster rendering. The explicit proxy representations directly represent continuous neural features in a discrete data structure, such as triplane Chan et al. (2022); Hu et al. (2023), voxel grid Fridovich-Keil et al. (2022); Sun et al. (2022), hash table Müller et al. (2022), or point sets Xu et al. (2022b). Recently, the 3D Gaussian Splatting methods Kerbl et al. (2023); Xu et al. (2023); Wu et al. (2023a); Cotton and Peyton (2024); Li et al. (2024) utilize 3D Gaussians as their underlying representation and adaptively densify and prune them during fitting, which offers impressive reconstruction quality. The fully explicit representation also provides real-time rendering speed. However, the 3D Gaussians are unstructured representation, and require per-scene optimization to achieve photo-realistic quality. In contrast, our work proposes a structured representation termed GaussianCube for 3D generative tasks.
Image-based 3D reconstruction. Compared to per-scene optimization, image-based 3D reconstruction methods Tatarchenko et al. (2019); Li et al. (2009); Tulsiani et al. (2017); Yu et al. (2021) can directly reconstruct 3D assets given images without optimization. PixelNeRF Yu et al. (2021) leverages an image feature encoder to empower the generalizability of NeRF. Similarly, pixel-aligned Gaussian approaches Charatan et al. (2023); Szymanowicz et al. (2023); Tang et al. (2024) follow this idea to design feed-forward Gaussian reconstruction networks. LRM Hong et al. (2023); He and Wang (2023) shows that transformers can also be scaled up for 3D reconstruction with large-scale training data, which is followed by hybrid Gaussian-triplane methods Zou et al. (2023); Xu et al. (2024) within the LRM frameworks. However, the limited number of Gaussians and spatially unstructured property hinders these methods from achieving high-quality reconstruction, which also makes it hard to extend them to 3D generative modeling.
3D generation. Previous works of SDS-based optimization Poole et al. (2022); Tang et al. (2023b); Xu et al. (2022a); Wang et al. (2024); Sun et al. (2023); Cheng et al. (2023); Chen et al. (2024) distill 2D diffusion priors Rombach et al. (2022) to a 3D representation with the score functions. Despite the acceleration Tang et al. (2023a); Yi et al. (2023) achieved by replacing NeRF with 3D Gaussians, generating high-fidelity 3D Gaussians using these optimization-based methods still requires costly test-time optimization. 3D-aware GANs Chan et al. (2022); Gao et al. (2022); Chan et al. (2021); Gu et al. (2021); Niemeyer and Geiger (2021); Deng et al. (2022); Xiang et al. (2022) can generate view-dependent images by training on single image collections. Nevertheless, they fall short in modeling diverse objects with complex geometry variations. Many recent works Wang et al. (2023); Müller et al. (2023); Gupta et al. (2023); Tang et al. (2023c); Shue et al. (2023) apply diffusion models for 3D generation using structured proxy 3D representations such as hybrid triplane Wang et al. (2023); Shue et al. (2023) or voxels Müller et al. (2023); Tang et al. (2023c). However, they typically need a shared implicit feature decoder across different assets, which greatly limits the representation expressiveness. Also, the inherent computational cost from NeRF leads to slow rendering speed, making it unsuitable for efficient training and rendering. Building upon the strong capability and rendering efficiency of Gaussian Splatting Kerbl et al. (2023), we propose a spatially structured Gaussian representation, making it suitable for 3D generative modeling. A concurrent work of He et al. (2024) also investigated transforming 3DGS into a volumetric representation. Their method confines the Gaussians to voxel grids during fitting and incorporates a specialized desification strategy. In contrast, our method only restricts the total number of Gaussians, adhering to the original splitting strategy and allowing unrestricted spatial distribution. This preserves the representation power during fitting. The subsequent OT-based voxelization yields spatially coherent arrangement with minimal global offset cost and hence effectively eases the difficulty of generative modeling.
Figure 2: Overall framework. Our framework comprises two main stages of representation construction and 3D diffusion. In the representation construction stage, given multi-view renderings of a 3D asset, we perform densification-constrained fitting to obtain 3D Gaussians with constant numbers. Subsequently, the Gaussians are voxelized into GaussianCube via Optimal Transport. In the 3D diffusion stage, our 3D diffusion model is trained to generate GaussianCube from Gaussian noise.
@@ 3 Method
Following prior works, our framework comprises two primary stages: representation construction and diffusion modeling. In representation construction phase, we first apply a densification-constrained 3DGS fitting algorithm for each object to obtain a constant number of Gaussians. These Gaussians are then organized into a spatially structured representation via Optimal Transport between the positions of Gaussians and centers of a predefined voxel grid. For diffusion modeling, we train a 3D diffusion model to learn the distribution of GaussianCubes. The overall framework is illustrated in Figure 2. We will detail our designs for each stage subsequently.
@@ 3.1 Representation Construction
We expect the 3D representation to be both structured, expressive and efficient. Despite Gaussian Splatting (GS) offers superior expressiveness and efficiency against NeRFs, it fails to yield fixed-length representations across different 3D assets; nor does it organize the data in a spatially structured format. To address these limitations, we introduce GaussianCube, which effectively overcomes the unstructured nature of Gaussian Splatting, while retaining both expressiveness and efficiency.
Formally, a 3D asset is represented by a collection of 3D Gaussians as introduced in Gaussian Splatting Kerbl et al. (2023). The geometry of the $i$-th 3D Gaussian $\bm{g}_{i}$ is given by
@@
Notably, the adaptive control is one of the most essential steps during the fitting process in GS Kerbl et al. (2023). It dynamically clones Gaussians in under-reconstructed regions, splits Gaussians in over-reconstructed regions, and eliminates those with irregular dimensions. Although the adaptive control substantially improves the fitting quality, it can lead to a varying number of Gaussians for different objects. Furthermore, the Gaussians are stored without a predetermined spatial order, resulting in an absence of an organized spatial structure. These aspects pose significant challenges to 3D generative modeling. To overcome these obstacles, we first introduce our densification-constrained fitting strategy to obtain a fixed number of free Gaussians. Then, we systematically arrange the resulting Gaussians within a predefined voxel grid via Optimal Transport, thereby achieving a spatially structured GS representation.
@@
Densification-constrained fitting. Our approach begins with the aim of maintaining a constant number of Gaussians $\bm{g}\in\mathbb{R}^{N_{\text{max}}\times C}$ across different objects during the fitting. A naive approach might involve omitting the densification and pruning steps in the original GS. However, we argue that such simplifications significantly harm the fitting quality, with empirical evidence shown in Table 4. Instead, we propose to retain the pruning process while imposing a new constraint on the densification phase. Specifically, if the current iteration comprises $N_{\text{c}}$ Gaussians and $N_{d}$ Gaussians need to be densified, we introduce a measure to prevent exceeding the predefined maximum of $N_{\text{max}}$ Gaussians (with $N_{\text{max}}$ set to $32,768$ in this work). This is achieved by selecting $N_{\text{max}}-N_{\text{c}}$ Gaussians with the largest view-space positional gradients from the $N_{d}$ candidates for densification in cases where $N_{d}>N_{\text{max}}-N_{\text{c}}$. Otherwise, all $N_{d}$ Gaussians are subjected to densification as in the original GS. Additionally, instead of performing the cloning and splitting in the same densification steps, we opt to perform each alternately without influencing each other. Upon completion of the entire fitting process, we pad Gaussians with $\alpha=0$ to reach the target count of $N_{\text{max}}$ without affecting the rendering results. The detailed fitting procedure is shown in Figure 3 (a).
Gaussian voxelization via Optimal Transport. To further organize the obtained Gaussians into a spatially structured representation for 3D generative modeling, we propose to map the Gaussians to a predefined structured voxel grid $\bm{v}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$ where $N_{v}=\sqrt[3]{N_{\text{max}}}$. Intuitively, we aim to “move” each Gaussian into a voxel grid while preserving their geometric relations as much as possible. To this end, we formulate this as an Optimal Transport (OT) problem Villani et al. (2009); Burkard and Cela (1999) between the Gaussians’ spatial positions $\{\bm{\mu}_{i},i=1,\ldots,N_{\text{max}}\}$ and the voxel grid centers $\{\bm{x}_{j},j=1,\ldots,N_{\text{max}}\}$. Let $\mathbf{D}$ be a distance matrix with $\mathbf{D}_{ij}$ being the moving distance between $\bm{\mu}_{i}$ and $\bm{x}_{j}$, i.e., $\mathbf{D}_{ij}=\|\bm{\mu}_{i}-\bm{x}_{j}\|^{2}$. The transport plan is represented by a binary matrix $\mathbf{T}\in\mathbb{R}^{N_{\text{max}}\times N_{\text{max}}}$, and the optimal transport plan is given by:
The solution is a bijective transport plan $\mathbf{T}^{*}$ that minimizes the total transport distances. We employ the Jonker-Volgenant algorithm Jonker and Volgenant (1988) to solve the OT problem. We organize the Gaussians according to the solutions, with the $j$-th voxel grid encapsulating the feature vector of the corresponding Gaussian $\bm{\theta}_{k}=\{\bm{\mu}_{k}-\bm{x}_{j},\bm{s}_{k},\bm{q}_{k},\alpha_{k},\bm{c}_{k}\}\in\mathbb{R}^{C}$, where $k$ is determined by the optimal transport plan (i.e., $\mathbf{T}^{*}_{kj}=1$). Note that we substitute the original Gaussian positions with offsets of the current voxel center to reduce the solution space for diffusion modeling. As a result, our fitted Gaussians are systematically arranged within a voxel grid $\bm{v}$ and maintain the spatial coherence.
@@
Formally, the generation procedure can be formulated into the inversion of a discrete-time Markov forward process. During the forward phase, we gradually add noise to $\bm{y}_{0}\sim p(\bm{y})$ and obtain a sequence of increasingly noisy samples $\{\bm{y}_{t}|t\in[0,T]\}$ according to
$\bm{y}_{t}:=\alpha_{t}\bm{y}_{0}+\sigma_{t}\bm{\epsilon},$ (3)
@@
where $\bm{\epsilon}\in\mathcal{N}(\mathbf{0},\bm{I})$ represents the added Gaussian noise, and $\alpha_{t},\sigma_{t}$ constitute the noise schedule which determines the level of noise added to destruct the original data sample. As a result, $\bm{y}_{T}$ will finally reach isotropic Guassian noise after sufficient destruction steps. By reversing the above process, we are able to perform the generation process by gradually denoise the sample starting from pure Gaussian noise $\bm{y}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$ until reaching $\bm{y}_{0}$. Our diffusion model is trained to denoise $\bm{y}_{t}$ into $\bm{y}_{0}$ for each timestep $t$, facilitating both unconditional and class-conditioned generation.

### v3

# GaussianCube: A Structured and Explicit Radiance Representation for 3D Generative Modeling

@@ Abstract
We introduce a radiance representation that is both structured and fully explicit and thus greatly facilitates 3D generative modeling. Existing radiance representations either require an implicit feature decoder, which significantly degrades the modeling power of the representation, or are spatially unstructured, making them difficult to integrate with mainstream 3D diffusion methods. We derive GaussianCube by first using a novel densification-constrained Gaussian fitting algorithm, which yields high-accuracy fitting using a fixed number of free Gaussians, and then rearranging these Gaussians into a predefined voxel grid via Optimal Transport. Since GaussianCube is a structured grid representation, it allows us to use standard 3D U-Net as our backbone in diffusion modeling without elaborate designs. More importantly, the high-accuracy fitting of the Gaussians allows us to achieve a high-quality representation with orders of magnitude fewer parameters than previous structured representations for comparable quality, ranging from one to two orders of magnitude. The compactness of GaussianCube greatly eases the difficulty of 3D generative modeling. Extensive experiments conducted on unconditional and class-conditioned object generation, digital avatar creation, and text-to-3D synthesis all show that our model achieves state-of-the-art generation results both qualitatively and quantitatively, underscoring the potential of GaussianCube as a highly accurate and versatile radiance representation for 3D generative modeling. Project page: https://gaussiancube.github.io/.
11footnotetext: Interns at Microsoft Research Asia. Equal advising. Corresponding author.
@@ 1 Introduction
The field of 3D generative modeling [55, 36, 5, 53, 46, 8, 17] has witnessed remarkable growth, driven by advancements in generative modeling [23, 18, 38, 15, 66, 27]. Most of the prior works in this domain leverage variants of Neural Radiance Field (NeRF) [35, 8, 53, 37] as their underlying 3D representations, which typically consist of an explicit structured proxy representation and an implicit feature decoder. However, such hybrid NeRF variants exhibit degraded representation power, particularly when used for generative modeling where a single implicit feature decoder is shared across all objects. Additionally, the high computational complexity of volumetric rendering leads to both slow rendering speed and extensive memory costs.
Recently, the emergence of 3D Gaussian Splatting (GS) [28] has enabled improved reconstruction quality and real-time rendering capabilities [64, 33, 58]. The fully explicit nature of 3DGS eliminates the need for a shared implicit decoder, providing another key advantage over NeRFs. Although 3DGS has been widely studied in scene reconstruction tasks, its spatially unstructured nature presents a significant challenge when applied to mainstream generative modeling frameworks.
To overcome these barriers, we introduce GaussianCube – an innovative radiance representation that is both structured and fully explicit, with strong fitting capabilities (see Table 1 for comparisons with prior works). The proposed approach first ensures high-accuracy fitting with a predefined number of free Gaussians, and subsequently organizes these Gaussians into a structured voxel grid. Such an explicit grid-based structure permits the seamless application of standard 3D convolutional architectures, such as U-Net, thereby eliminating the need for complex, specialized network designs [69, 55] that are often necessary with unstructured or implicitly decoded representations.
@@
Structuring 3D Gaussians without sacrificing fitting quality is not a trivial task. A naive starting point would be obtaining a fixed number of Gaussians by omitting the densification and pruning steps in GS. However, such simplification fails to lead the Gaussians close to the object surfaces and results in significant quality degradation. In contrast, we propose a densification-constrained fitting strategy, which retains the original pruning process yet constrains the number of Gaussians that perform densification, ensuring the total does not exceed a predefined maximum $N^{3}$. For the subsequent structuralization, we allocate the Gaussians across an $N\times N\times N$ voxel grid using Optimal Transport (OT). Consequently, our fitted Gaussians are systematically arranged within the voxel grid, with each voxel containing the features of a Gaussian. The proposed OT-based structuralization achieves maximal spatial correspondence, characterized by minimal total transport distances, while preserving the expressive power of 3DGS.
[…]
Instant-NGP [35] 26.63×26.63$\times$26.63 ×
Neural Voxels [53] 145.9×145.9$\times$145.9 ×
Triplane [8] 13.7×13.7$\times$13.7 ×
Gaussian Splatting [28] 4.0×4.0$\times$4.0 ×
Our GaussianCube $\mathbf{1.0\times}$
Table 1: Comparison with previous 3D representations with respect to spatial structure, explicitness, real-time rendering capability, and relative parameter count (Rel. Parameters) for representations of comparable quality.
The structured nature of GaussianCube enables us to perform efficient 3D diffusion [23] modeling for the following three reasons: 1) It allows the use of standard 3D U-Net as our backbone for diffusion modeling without elaborate designs. 2) The spatial coherence of GaussianCube permits the use of standard 3D convolutions to capture the correlations among neighboring Gaussians, facilitating efficient feature extraction. 3) GaussianCube enables high-quality fitting with orders of magnitude fewer parameters than prior structured representations of similar quality. Since recent works [30, 3] have demonstrated diffusion models’ struggle in handling high-dimensional distributions, the compactness of GaussianCube significantly reduces the modeling difficulty of the generative framework.
@@
We conduct comprehensive experiments to verify the efficacy of our approach. The model’s capability for unconditional and class-conditioned generation is evaluated on the ShapeNet [9] and OmniObject3D [59] datasets. Both the quantitative and qualitative comparisons indicate that our model surpasses all previous methods. We also perform digital avatar generation on a synthetic avatar dataset [57]. Our model is capable of producing high-fidelity 3D avatars conditioned on single portrait images, excelling beyond prior art in both identity preservation and detail creation. Additionally, we assess our model’s capacity for the challenging text-to-3D creation task on Objaverse [12]. Our model demonstrates competitive performance both quantitatively and qualitatively, producing results consistent with the given text prompts in just 5 seconds. All experiments show the strong capabilities of our GaussianCube and suggest its potential as a powerful and versatile 3D representation for a variety of applications. Some generated samples of our method are presented in Figure 1.
Figure 1: Our diffusion model is able to create diverse objects with complex geometry and rich texture details (top three rows). Our method also supports creating high-fidelity digital avatars (the forth row) conditioned on single portrait images (visualized in dashed boxes) and high-quality 3D assets given text prompts (the fifth row).
@@ 2 Related Work
Radiance field representation. Radiance fields model ray interactions with scene surfaces and can be in either implicit or explicit forms. Early works of neural radiance fields (NeRFs) [35, 67, 40, 1, 42] are often in an implicit form, which represents scenes without defining geometry. These works optimize a continuous scene representation using volumetric ray-marching that leads to extremely high computational costs. Recent works introduce the use of explicit proxy representation [8, 24, 16, 48, 37, 63] followed by an implicit feature decoder to enable faster rendering. Recently, the 3D Gaussian Splatting methods [28, 64, 58, 11, 29] utilize 3D Gaussians as their underlying representation and offer impressive reconstruction quality. The fully explicit representation also provides real-time rendering speed. However, the 3D Gaussians are unstructured representation, and require per-scene optimization to achieve photo-realistic quality. In contrast, our work proposes a structured representation termed GaussianCube for 3D generative tasks.
3D generation. Previous works of SDS-based optimization [41, 52, 62, 49, 10, 50, 65] distill 2D diffusion priors [44] to a 3D representation with the score functions, but these works are notably time-intensive, often taking several minutes to hours. While 3D-aware GANs [8, 17, 7, 19, 39, 14, 61] facilitate view-dependent image generation from single-image collections, they struggle to capture the complexity of diverse objects with intricate geometric variations [60]. Although recent works[55, 36, 20, 53, 46] have utilized diffusion models with structured proxy representations for 3D generation, the use of a shared implicit feature decoder across different assets restricts expressiveness and the computational demands of NeRF hinder efficient training. In contrast, we introduce a structured and fully explicit radiance representation for 3D generative modeling, building upon 3DGS [28]. A concurrent work of [21] includes elaborate designs to form the Gaussians into volumetric representation during fitting, yet does not thoroughly address global correspondence. In contrast, our approach only restricts the total count of Gaussians while allowing freedom in their spatial distribution during the fitting. We then organize these Gaussians into a voxel grid using Optimal Transport, which yields a spatially coherent arrangement with minimal global offset cost, effectively easing the difficulty of generative modeling.
@@ 3 Method
Following prior works, our framework comprises two primary stages as shown in Figure 2: representation construction and diffusion modeling. In representation construction phase, we first apply a densification-constrained 3DGS fitting algorithm for each object to obtain a constant number of Gaussians. These Gaussians are then organized into the proposed spatially structured GaussianCube via Optimal Transport between the positions of Gaussians and centers of a predefined voxel grid. For diffusion modeling, we train a 3D diffusion model to learn the distribution of GaussianCubes. We will detail our designs for each stage subsequently.
Figure 2: Overall framework. Our framework comprises two main stages of representation construction and 3D diffusion. In the representation construction stage, given multi-view renderings of a 3D asset, we perform densification-constrained fitting to obtain 3D Gaussians with constant numbers. Subsequently, the Gaussians are structured into GaussianCube via Optimal Transport. In the 3D diffusion stage, our 3D diffusion model is trained to generate GaussianCube from Gaussian noise.
@@ 3.1 Representation Construction
We build our GaussianCube upon 3DGS, an explicit representation that offers impressive fitting quality and real-time rendering speed. However, it fails to yield Gaussians of fixed length since the adaptive density control during GS fitting can lead to a varying number of Gaussians for different objects. Furthermore, the lack of a predetermined spatial ordering for Gaussians leads to a disorganized spatial structure. These aspects pose significant challenges to 3D generative modeling. To overcome these obstacles, we first introduce our densification-constrained fitting strategy to obtain a fixed number of free Gaussians. Then, we systematically arrange the resulting Gaussians within a predefined voxel grid via Optimal Transport, thereby achieving a structured and explicit radiance representation.
Formally, a 3D asset is represented by a collection of 3D Gaussians as introduced in Gaussian Splatting [28]. The geometry of the $i$-th 3D Gaussian $\bm{g}_{i}$ is given by
@@
Densification-constrained fitting. Our approach begins with the aim of maintaining a constant number of Gaussians $\bm{g}\in\mathbb{R}^{N_{\text{max}}\times C}$ across different objects during the fitting. A simplistic approach might involve omitting the densification and pruning in the original GS. However, we argue that such simplifications significantly harm the fitting quality, with empirical evidence shown in Table 6. Instead, we propose to retain the pruning process while imposing a new constraint on the densification phase as shown in Figure 3 (a). The fitting process encompasses several distinct stages: 1) Densification Detection: Assuming the current iteration includes $N_{\text{c}}$ Gaussians, we identify densification candidates by selecting those with view-space position gradient magnitudes exceeding a predefined threshold $\tau$. We denote the number of candidates as $N_{d}$. 2) Candidate sampling: To prevent exceeding the predefined maximum of $N_{\text{max}}$ Gaussians, we select $\min{(N_{\text{max}}-N_{\text{c}},N_{d})}$ Gaussians with the largest view-space positional gradients from the candidates for densification. 3) Densification: We modify the densification approach by alternating between cloning and splitting actions into separate steps. 4) Pruning Detection and Pruning: We identify and remove the Gaussians with $\alpha$ less than a small threshold $\epsilon$. After completing the fitting process, we pad Gaussians with $\alpha=0$ to reach the target count of $N_{\text{max}}$ without affecting the rendering results. Benefiting from our proposed strategy, we attain a high-quality representation with orders of magnitude fewer parameters compared to existing works of similar quality, which significantly reduces the modeling difficulty for the diffusion models.
Gaussian structuralization via Optimal Transport. To further organize the obtained Gaussians into a spatially structured representation for 3D generative modeling, we propose to map the Gaussians to a predefined structured voxel grid $\bm{v}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$ where $N_{v}=\sqrt[3]{N_{\text{max}}}$. Intuitively, we aim to “move” each Gaussian into a voxel while preserving their geometric relations as much as possible. While naive approaches such as nearest neighbor transport fall short in conserving these relations due to disregard for global arrangement with evidence shown in Figure 10, we formulate this as an Optimal Transport (OT) problem [54, 4] between the Gaussians’ spatial positions $\{\bm{\mu}_{i},i=1,\ldots,N_{\text{max}}\}$ and the voxel grid centers $\{\bm{x}_{j},j=1,\ldots,N_{\text{max}}\}$. Let $\mathbf{D}$ be a distance matrix with $\mathbf{D}_{ij}$ being the moving distance between $\bm{\mu}_{i}$ and $\bm{x}_{j}$, i.e., $\mathbf{D}_{ij}=\|\bm{\mu}_{i}-\bm{x}_{j}\|^{2}$. The transport plan is represented by a binary matrix $\mathbf{T}\in\mathbb{R}^{N_{\text{max}}\times N_{\text{max}}}$, and the optimal transport plan is given by:
The solution is a bijective transport plan $\mathbf{T}^{*}$ that minimizes the total transport distances. We employ the Jonker-Volgenant algorithm [25] to solve the OT problem. We provide a 2D illustration in Figure 3 (b). We organize the Gaussians according to the solutions, with the $j$-th voxel encapsulating the feature vector of the corresponding Gaussian $\bm{\theta}_{k}=\{\bm{\mu}_{k}-\bm{x}_{j},\bm{s}_{k},\bm{q}_{k},\alpha_{k},\bm{c}_{k}\}\in\mathbb{R}^{C}$, where $k$ is determined by the optimal transport plan (i.e., $\mathbf{T}^{*}_{kj}=1$). Note that we replace the original Gaussian positions with offsets of the current voxel center to reduce the solution space for diffusion models. As a result, our fitted Gaussians are systematically arranged within a voxel grid $\bm{v}$ and preserve the spatial correspondence of neighboring Gaussians, which further facilitates generative modeling.
@@
Formally, the generation procedure can be formulated into the inversion of a discrete-time Markov forward process. During the forward phase, we gradually add noise to $\bm{y}_{0}\sim p(\bm{y})$ and obtain a sequence of increasingly noisy samples $\{\bm{y}_{t}|t\in[0,T]\}$ according to $\bm{y}_{t}:=\alpha_{t}\bm{y}_{0}+\sigma_{t}\bm{\epsilon}$, where $\bm{\epsilon}\in\mathcal{N}(\mathbf{0},\bm{I})$ represents the added Gaussian noise, and $\alpha_{t},\sigma_{t}$ constitute the noise schedule. As a result, $\bm{y}_{T}$ will finally reach isotropic Gaussian noise after sufficient destruction steps. By reversing the above process, we are able to perform the generation process by gradually denoise the sample starting from pure Gaussian noise $\bm{y}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$ until reaching $\bm{y}_{0}$. Our diffusion model is trained to denoise $\bm{y}_{t}$ into $\bm{y}_{0}$ for each timestep $t$, facilitating both unconditional and conditional generation.